OpenStudy (anonymous):

factor : 25 - 100y^2

OpenStudy (anonymous):

= (5 + 10y)(5 - 10y)

OpenStudy (anonymous):

Rebecca1233 - By now, you should know the difference of squares. Takes less time to learn and do than typing all these questions. That is just my opinion.

OpenStudy (anonymous):

I can do them i just get confused sometime so if someone can give me the answer then i can compare and etc sometimes i need help.. i thank these people who help me!

OpenStudy (muhammad_nauman_umair):

25 - 100y^2= (5 + 10y)(5 - 10y)

OpenStudy (precal):

@rebecca1233 we don't mind helping you understand the process. You do yourself no favors by not responding to us when we ask you a question. We are tryng to help you learn so that you will do better on an upcoming exam or future math course.